Operations Report Oct 24th

  • Crew 197 Operations Report 24Oct18
    SOL: 3
    Name of person filing report: Jim Ehrhart

Non-nominal systems: solar/batteries
Notes on non-nominal systems: batteries apparently on site, to be
installed Friday
Generator (hours run): 23.5, generator on almost full-time, down 5 min
at approx. 5:30pm to add coolant
Solar— SOC % (Before generator is run at night): N/A
Diesel: 75%
Propane – psi: 45
Ethanol Free Gasoline (5 Gallon containers for ATV) – gallons: Not
checked, ATV not in use

Water (trailer) – gallons: N/A trailer not in use
Water (static) – gallons: estimated 90% full
Trailer to Static Pump used – no or yes: No
Water (loft) – Static to Loft Pump used – yes or no: Yes
Water Meter: morning fill, from 10 to 50, meter 138088.98; evening
fill from 10 to 50, meter 138112.92
Toilet tank emptied: no or yes: Yes

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): No
Oil Added? No
ATV Fuel Used: Gals No
# Hours the ATVs were Used today: 0
Notes on ATVs: N/A

Deimos rover used: No, out of service
Hours: 134
Beginning charge: N/A
Ending charge: N/A
Currently charging: No
Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: Unk
Beginning charge: Unk
Ending charge: Unk
Currently charging: Unk
Spirit rover used: Yes
Hours: 42.3
Beginning charge: 100
Ending charge: 99
Currently charging: Yes
Opportunity rover used: Yes
Hours: 36.1 hrs
Beginning charge: 100
Ending charge: 100
Currently charging: No
Curiosity rover used: Yes
Hours: 55.3 hrs
Beginning charge: 100
Ending charge: 97
Currently charging: Yes

HabCar used and why, where? No

General notes and comments:

Summary of internet: Internet connectivity and bandwidth fine
Summary of suits and radios: All suits performed nominally. One
charger missing and one wonky – will create two new chargers from
existing spare parts tomorrow with suit engineers. Suits experienced
wear and tear on EVAs, some quilting torn in rover ingress/egress, two
scratches on one helmet
All radios reviewed and settings standardized. 13 units in good shape
and fully charged. 4 units not in good condition and set aside as
Summary of Hab operations: All systems nominal. Fan/heater air filter
replaced. Loft water filter replaced.
Summary of GreenHab operations: Nominal, smoke detector batteries
replaced and tested
Summary of ScienceDome operations: N/A, not in use due to battery issue
Summary of RAMM operations: Nominal, very disorderly, needs cleaning
and organization
Summary of health and safety issues: Health and Safety Officer perform
checks – see report.
Questions, concerns and requests to Mission Support:
– Battery meter needed to check random spares found in RAMM.
– Suggest light in loft

Copyright © The Mars Society. All rights reserved. | Main Site